OpenStudy (anonymous):

Can someone help me with this please? I don't want the answers, I want someone to help me understand how to solve the question. A function is shown below: f(x) = x3 + 2x2 - x - 2 Part A: What are the factors of f(x)? Show your work. Part B: What are the zeros of f(x)? Show your work. Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x). Thanks!

OpenStudy (anonymous):

@zepdrix @mathmale @mathstudent55 @triciaal @ganeshie8 @Directrix

zepdrix (zepdrix):

Figure out part A? :) Looks like grouping again.

OpenStudy (anonymous):

Yeah, I was thinking of doing that later. I thot I took it out of the question... Eh I'll just do it here

zepdrix (zepdrix):

Well you need part A for the rest of the problem XD so make sure you can do that

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

f(x) = (x3 + 2x2 - x - 2) First group: f(x) = (x3 + 2x2) - (x - 2) Find GCF First Binomial x2 Second Binomial -1 Factor out: f(x) = x2(x + 2) -1(x + 2) (x2 - 1)(x + 2) Is this correct @zepdrix I think I might have done something wrong...

zepdrix (zepdrix):

Looks good so far :) We can go a little bit further. Here is something to think about:

OpenStudy (anonymous):

wait wait wait wait

zepdrix (zepdrix):

No wait >:U\[\large\rm x^2-1\quad=\quad x^2-1^2\]

zepdrix (zepdrix):

See your next step? :)

OpenStudy (anonymous):

but theres something incorrect f(x) = x2(x + 2) -1(x + 2) Factoring out -1 would give f(x) = x2(x + 2) -1(-x + 2) Not f(x) = x2(x + 2) -1(x + 2)

OpenStudy (anonymous):

D:

zepdrix (zepdrix):

oh i missed that boo boo. You should've grouped like this:

zepdrix (zepdrix):

\[\large\rm (x^3+2x^2)+(-x-2)\]You have to input a plus sign to put the brackets the way you would like. You can't just drop brackets like this:\[\large\rm (x^3+2x^2)-(x-2)\]I don't think that's what you did, just a note.

zepdrix (zepdrix):

\[\large\rm x^3+2x^2-x-2\]\[\large\rm (x^3+2x^2)+(-x-2)\]\[\large\rm x^2(x+2)-1(x+2)\]\[\large\rm (x^2-1)(x+2)\]These are the correct steps. Confused by any of this?

OpenStudy (anonymous):

@zepdrix Alright, I get it so far. It makes sense.

OpenStudy (anonymous):

So for A, the factors are (x2−1)(x+2) correct @zepdrix ???

zepdrix (zepdrix):

No, we would like `linear factors`, so we factor further.

OpenStudy (anonymous):

0-0

zepdrix (zepdrix):

\[\large\rm x^2-1\quad=\quad x^2-1^2\]Remember your `difference of squares` formula?

zepdrix (zepdrix):

\[\large\rm a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

so how would i use that in this???

OpenStudy (anonymous):

btw thx zep, ur awesome. :)

OpenStudy (anonymous):

@zepdrix hmmm...x^2−1=x2−1^2. I don't really get how you suddenly got 1^2 oh 1 * 1 is still 1 nvm. Alright that makes sesnse so far. Equation is a2−b2=(a−b)(a+b) Im gonna go on a limb here and say a2 is x2 and b2 is 1^2. So x^2 - 1^2 = (x - 1)(x + 1) Therefore, the fully factored form would be (x - 1)(x + 1) Correct @zepdrix ?

OpenStudy (anonymous):

@zepdrix So the factors of f(x) would be (x - 1)(x + 1), correct?

zepdrix (zepdrix):

Good good good. That gives us our fully factored expression for part A: (x-1)(x+1)(x+2)

OpenStudy (anonymous):

Alright, the fully factored form is (x-1)(x+1)(x+2) for Part A?

zepdrix (zepdrix):

\[\large\rm (x^2-1)(x+2)\]\[\large\rm (x-1)(x+1)(x+2)\]Yes.

OpenStudy (anonymous):

got it, thanks!

OpenStudy (anonymous):

Alright, so to calculate the x intercepts you have to make the equation equal to 0, so i think to calculate the 0's you have to make x = 0 first, and then y = 0? Or is that not true...

zepdrix (zepdrix):

\[\large\rm y=(x-1)(x+1)(x+2)\]Looking for zeros, or x-intercepts,\[\large\rm 0=(x-1)(x+1)(x+2)\]Applying your `Zero-Factor Property`:\[\large\rm 0=(x-1),\qquad\qquad\qquad 0=(x+1),\qquad\qquad\qquad 0=(x+2)\]And solve for x in each case.

zepdrix (zepdrix):

Setting x=0 is a separate issue. That gives us the `y-intercept`. it doesn't appear that they want information though, so we don't need to do that.

OpenStudy (anonymous):

Alright, so the 0s are just the x intercepts, correct @zepdrix ?

zepdrix (zepdrix):

Yes, these words mean all the same thing: Zeros, x-intercepts, solutions. They'll use those three different words a lot, so get comfortable with the fact that they mean the same thing.

OpenStudy (anonymous):

\[0=(x−1) 0=(x+1) 0=(x+2)\] Alright In the first equation x = 1 Second equation x = -1 Third equation x = -2

zepdrix (zepdrix):

Good :) You've got your zeroes.

OpenStudy (anonymous):

HOORAY! On to part C

OpenStudy (anonymous):

Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x).

zepdrix (zepdrix):

To graph it? You would label your x-intercepts to start.

OpenStudy (anonymous):

f(x) = x3 + 2x2 - x - 2 Right right, that is a good start lol

OpenStudy (anonymous):

Wait, this has 3 intercepts, does that mean 2 vertexes?

zepdrix (zepdrix):

Yes, and without the use of calculus, we can't graph those very easily. So I don't think they want you to worry about that too much.

OpenStudy (anonymous):

Hold on for a minute please

zepdrix (zepdrix):

\(\large\rm f(x) = x^3 + 2x^2 - x - 2\) If you let x=0, you have: \(\large\rm f(x) = 0 + 0 - 0 - 2\) So you can also graph your y-intercept if you like. y=-2

OpenStudy (anonymous):

@zepdrix Alright I'm back. I'm fairly sure I somehow found the vertex in my other question tho

OpenStudy (anonymous):

Forgot how...

OpenStudy (anonymous):

@zepdrix Dont I need to plot all the intercepts to be able to graph f(x)? Cause then I can just draw a line in between, right?

zepdrix (zepdrix):

No, not in this type of problem you didn't. The vertices aren't even rational numbers for this one. :) Really weird decimal values.

zepdrix (zepdrix):

Yes, you can connect things somewhat effectively after you do that.

zepdrix (zepdrix):

But you also need to know the `end behavior` of the function.

OpenStudy (anonymous):

but how would i graph without the vertexes?

zepdrix (zepdrix):

You would just sort of... "guess" how high the vertices go. You wouldn't be able to do the same thing you did with a parabola, vertex form is nice and easy with a parabola :)) not with this bad boy.

zepdrix (zepdrix):

Sec, lemme get some graph paper for us.

zepdrix (zepdrix):



OpenStudy (anonymous):

@zepdrix Sorry, bathroom break. Alright I see the 0

OpenStudy (anonymous):

@zepdrix Did you get banned?

zepdrix (zepdrix):

Lol, what? Why would I ever get banned? XD

OpenStudy (anonymous):

I don't know, conversing with an OS fugitive whose account has been banned?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Alright, I still don't understand how you would be able to graph this without knowing the vertexes @zepdrix .

zepdrix (zepdrix):

So if we graph the intercepts, that gives us a basic idea.